OpenStudy (anonymous):

Find the exact value of the real number y. y = csc^-1(1)

OpenStudy (anonymous):

i believe it is pi/4

OpenStudy (anonymous):

csc is the inverse of sine.

OpenStudy (e.mccormick):

Not quite...

OpenStudy (anonymous):

so its basically asking you to find get angle when sin is (1)^-1.

OpenStudy (e.mccormick):

(not quite \(\frac{\pi}{4}\) that is...)

OpenStudy (anonymous):

ok, i tried it again, and i got pi/2 this time, i know it isnt 2pi because that would be a full circle, and it isnt pi because that would be half a circle

OpenStudy (e.mccormick):

Yah, \(\frac{\pi}{2}\). The top of the unit circle.

OpenStudy (e.mccormick):

Have you seen the graphs of the periodic functions? It might help you understand the concept a little better.

OpenStudy (e.mccormick):

If you look at this graph: https://www.desmos.com/calculator/xr6ycnqkfr You can turn off the parts you do not need. Leave on the line \(x=\tfrac{\pi}{2}\) and \(\csc x\). You can see that at that value of x, the y value is 1. Well, \(\csc^{-1}\) or arccsc is the inverse function, so that question meant "When Y is 1, what is the X value for csc?" Hope that helps you see it another way.